# LeWiDi-2025 at NLPerspectives: The Third Edition of the Learning with Disagreements Shared Task

### v1

@@ Abstract
Many researchers have reached the conclusion that ai models should be trained to be aware of the possibility of variation and disagreement in human judgments, and evaluated as per their ability to recognize such variation. The LeWiDi series of shared tasks on Learning With Dis- agreements was established to promote this approach to training and evaluating ai models, by making suitable datasets more accessible and by developing evaluation methods. The third edition of the task builds on this goal by ex- tending the LeWiDi benchmark to four datasets spanning paraphrase identification, irony detec- tion, sarcasm detection, and natural language inference, with labeling schemes that include not only categorical judgments as in previous editions, but ordinal judgments as well. Another novelty is that we adopt two complementary paradigms to evaluate disagreement-aware systems: the soft-label approach, in which models predict population-level distributions of judgments, and the perspectivist approach, in which models predict the interpretations of individual annotators. Crucially, we moved beyond standard metrics such as cross-entropy, and tested new evaluation metrics for the two paradigms. The task attracted diverse participation, and the results provide insights into the strengths and limitations of methods to modeling variation. Together, these contributions strengthen LeWiDi as a framework and provide new resources, benchmarks, and findings to support the development of disagreement-aware technologies.
LeWiDi-2025 at NLPerspectives: The Third Edition of the Learning with Disagreements Shared Task
Elisa Leonardelli1, Silvia Casola2, Siyao Peng2, Giulia Rizzi3, Valerio Basile4, Elisabetta Fersini4, Diego Frassinelli 2, Hyewon Jang5, Maja Pavlovic6 Barbara Plank2, Massimo Poesio6,7 1Fondazione Bruno Kessler, 2LMU Munich & MCML, 3Università Milano Bicocca, 4Università di Torino, 5University of Gothenburg, 6Queen Mary University of London, 7Utrecht University
@@ 2 The LeWiDi 3 Benchmark
The four selected datasets are summarized in Table 2, illustrated with examples in Table 2, and described in detail in the following sections.
@@ 2.2 The MultiPICo dataset (MP)
The MP dataset Casola et al. (2024) is a multilingual perspectivist corpus consisting of short exchanges from Twitter and Reddit. Each entry in the corpus represents a post-reply pair. Crowdsourced workers had to determine whether the reply was ironic given the post (binary label). The corpus includes 11 languages: Arabic, Dutch, English, French, German, Hindi, Italian, Portuguese, and Spanish. It also contains sociodemographic information about the annotators, including gender, age, nationality, race, and student or employment status. While the statistics may vary slightly across languages, each post-reply pair is typically annotated by an average of 5 workers.
@@ 2.4 The Paraphrase Detection dataset (Par)
Dataset
(detection of)
Example
Annotations
(Task B)
AnnotatorId:Label
Soft labels
(Task A)
Label:Probability
CSC
context: "You walk into the room and
Steve is there and Steve says ‘‘hi!’’"
response: "hi"
A812:1, A813:3,
A814:1, A815:2
[1:0.5, 2:0.25, 3:0.25,
4:0, 5:0, 6:0]
MP
post: "@USER Oh dear"
reply: "@USER It’s ok, wine has fixed
everything"
A26:1, A64:1,
A70:1
[0:0, 1:1]
Par
Q1: "Have you seen an alien craft?"
Q2: "Have you ever seen an alien?"
A1:-1, A2:-3,
A3:5, A4:4
[-5:0, -4:0, -3:0.25, -2:0,
-1:0.25, 0:0, 1:0, 2:0, 3:0,
4:0.25, 5:0.25]
VEN
context: "yeah i can believe that"
statement: "I agree with what you said."
A1:E, A2:N,
A3:N, A4:E
[C:[0:1, 1:0]
E:[0:0.5, 1:0.5]
N:[0:0.5, 1:0.5]]
@@ 5.1 Systems overview
Opt-ICL Sorensen and Choi (2025) combines in-context learning (ICL) with fine-tuning in a two-stage approach. They first apply post-training, by exposing an LLM to over 40 datasets rich in human disagreement Sorensen et al. (2025), and then, for each dataset, conduct supervised fine-tuning, using in-context demonstrations from all the individual annotators along with annotator demographics. At inference, the model performs per-rater prediction by constructing a prompt with as many training examples from that annotator as possible, followed by the input to be labeled. They derive soft label distributions from perspectivist predictions.
@@
DeMeVa Ignatev et al. (2025) employs LLMs with ICL, modeling perspectivism through annotators’ past behavior. They focus on criteria for selecting demonstrative examples for LLMs (10 per annotator), comparing semantic and label-based strategies, with the latter performing better for multi-label datasets. They derive soft label distributions from perspectivist predictions.
@@
twinhter Nguyen and Van Thin (2025) built a BERT-based model that integrates annotator perspectives by creating a new (text, annotator) pair. They create a separate training instance for each annotator’s view and combine it with their background information when available, enabling the model to capture individual interpretations of the same input.
@@
McMaster Sanghani et al. (2025) implemented a demographic-aware RoBERTa model that incorporates information such as age, gender, nationality, and evaluated it across all four datasets. The authors find that nationality and ethnicity in particular show the largest gains in performance, while also noting the limitations of relying on such features.
@@
BoN Appetite Team Ruiz et al. (2025) investigated three test-time scaling methods, a way to improve LLMs performances: two benchmark algorithms (Model Averaging and Majority Voting), and a Best-of-N (BoN) sampling method. Their results show that the benchmark methods (Averaging and Voting) reliably boost performance, while BoN sampling does not transfer well from mathematical domains.
@@
PromotionGo (Huang et al., 2025) submitted only to the MP-Task A with an XLM-R–based system, ranking first. They deployed three main strategies to develop a competitive system: data augmentation, including lexical swaps, prompt-based reformulation, and large-scale back-translation into nine languages; optimization for alignment to the evaluation metric (Manhattan Distance) by using L1 loss as a loss function; ensemble learning, by training multiple models on shuffled data splits and averaging predictions to improve robustness.
@@
Uncertain Mis(Takes) Staliūnaitė and Vlachos (2025) addressed only the VEN-Task A, ranking first. They aim to quantify ambiguity in NLI instances, relying on the hypothesis that if a given instance is ambiguous, then the explanations for different labels will not entail one another. For each item, they generate 128 LLM explanations. With a fine-tuned entailment model they cluster them and quantify their Semantic Entropy (SE). The explanation clusters’ SE scores are combined with text embeddings for soft label distribution prediction.
@@
NLP-ResTEAM Sarumi et al. (2025) proposed a multi-task architecture. Special ‘tokens’ are added to the input, including several tokens aiming at modeling the annotators based on their ID, their demographic features, their annotation behavior, or combinations of those. The system produces two outputs from a textual input and an annotator’s information: one is a soft-label, the other a prediction of that specific annotator’s (hard) label.
@@
LPI-RIT Sawkar et al. (2025) builds upon the DisCo (Distribution from Context) architecture Weerasooriya et al. (2023), a neural model that jointly predicts item-level, annotator-level, and per-annotator label distributions. They tackled both soft-label and perspectivist tasks simultaneously. They also introduced several extensions to DisCo, such as integrating annotator metadata through pretrained sentence encoders, and modified loss functions to better align with evaluation metrics.
@@ 6.1 Results and statistics
Table 3 and 4 report the overall leaderboard for Task A and Task B respectively. If a team did not submit predictions for a particular dataset or task, we used the random baseline results to compute the overall ranks and average positions. Ranks were calculated with statistical ties taken into account. Specifically, we used the Wilcoxon signed-rank test at the instance level to identify clusters of tied systems. Predictions that were not significantly different (p = 0.05) from the top-performing system in a given cluster were considered ties. A new cluster was formed when a system’s performance was found to be statistically different from that of the best-performing system in the previous cluster. Leadboards for each specific dataset are reported in Appendix B.

### v2

# LeWiDi-2025 at NLPerspectives: Third Edition of the Learning with Disagreements Shared Task

@@ Abstract
Many researchers have reached the conclusion that ai models should be trained to be aware of the possibility of variation and disagreement in human judgments, and evaluated as per their ability to recognize such variation. The LeWiDi series of shared tasks on Learning With Dis- agreements was established to promote this approach to training and evaluating ai models, by making suitable datasets more accessible and by developing evaluation methods. The third edition of the task builds on this goal by extending the LeWiDi benchmark to four datasets spanning paraphrase identification, irony detection, sarcasm detection, and natural language inference, with labeling schemes that include not only categorical judgments as in previous editions, but ordinal judgments as well. Another novelty is that we adopt two complementary paradigms to evaluate disagreement-aware systems: the soft-label approach, in which models predict population-level distributions of judgments, and the perspectivist approach, in which models predict the interpretations of individual annotators. Crucially, we moved beyond standard metrics such as cross-entropy, and tested new evaluation metrics for the two paradigms. The task attracted diverse participation, and the results provide insights into the strengths and limitations of methods to modeling variation. Together, these contributions strengthen LeWiDi as a framework and provide new resources, benchmarks, and findings to support the development of disagreement-aware technologies.
@@ 2 The LeWiDi 3 Benchmark
The four selected datasets are summarized in Table 2, illustrated with examples in Table 2.4, and described in detail in the following sections.
@@ 2.2 The MultiPICo dataset (MP)
The MP dataset Casola et al. (2024) is a multilingual perspectivist corpus consisting of short exchanges from Twitter and Reddit. Each entry in the corpus represents a post-reply pair. Crowdsourced workers had to determine whether the reply was ironic given the post (binary label). The corpus includes 9 languages: Arabic, Dutch, English, French, German, Hindi, Italian, Portuguese, and Spanish. It also contains sociodemographic information about the annotators, including gender, age, nationality, race, and student or employment status. While the statistics may vary slightly across languages, each post-reply pair is typically annotated by an average of 5 workers.
@@ 2.4 The Paraphrase Detection dataset (Par)
Dataset (detection of) Example Annotations (Task B) AnnotatorId:Label Soft labels (Task A) Label:Probability
CSC (Sarcasm) context: "You walk into the room and Steve is there and Steve says ‘‘hi!’’" response: "hi" A812:1, A813:3, A814:1, A815:2 [1:0.5, 2:0.25, 3:0.25, 4:0, 5:0, 6:0]
MP (Irony) post: "@USER Oh dear" reply: "@USER It’s ok, wine has fixed everything" A26:1, A64:1, A70:1 [0:0, 1:1]
Par (Paraphrase) Q1: "Have you seen an alien craft?" Q2: "Have you ever seen an alien?" A1:-1, A2:-3, A3:5, A4:4 [-5:0, -4:0, -3:0.25, -2:0, -1:0.25, 0:0, 1:0, 2:0, 3:0, 4:0.25, 5:0.25]
VEN (NLI) context: "yeah i can believe that" statement: "I agree with what you said." A1:E, A2:N, A3:N, A4:E [C:[0:1, 1:0] E:[0:0.5, 1:0.5] N:[0:0.5, 1:0.5]]
@@ 5.1 Systems overview
Opt-ICL Sorensen,Taylor and Choi,Yejin (2025) combines in-context learning (ICL) with fine-tuning in a two-stage approach. They first apply post-training, by exposing an LLM to over 40 datasets rich in human disagreement Sorensen et al. (2025), and then, for each dataset, conduct supervised fine-tuning, using in-context demonstrations from all the individual annotators along with annotator demographics. At inference, the model performs per-rater prediction by constructing a prompt with as many training examples from that annotator as possible, followed by the input to be labeled. They derive soft label distributions from perspectivist predictions.
@@
DeMeVa Ignatev,Daniil et al. (2025) employs LLMs with ICL, modeling perspectivism through annotators’ past behavior. They focus on criteria for selecting demonstrative examples for LLMs (10 per annotator), comparing semantic and label-based strategies, with the latter performing better for multi-label datasets. They derive soft label distributions from perspectivist predictions.
@@
twinhter Dang and Thin,Dang (2025) built a BERT-based model that integrates annotator perspectives by creating a new (text, annotator) pair. They create a separate training instance for each annotator’s view and combine it with their background information when available, enabling the model to capture individual interpretations of the same input.
@@
McMaster Sanghani,Aadi et al. (2025) implemented a demographic-aware RoBERTa model that incorporates information such as age, gender, nationality, and evaluated it across all four datasets. The authors find that nationality and ethnicity in particular show the largest gains in performance, while also noting the limitations of relying on such features.
@@
BoN Appetite Team Ruiz,Tomas et al. (2025) investigated three test-time scaling methods, a way to improve LLMs performances: two benchmark algorithms (Model Averaging and Majority Voting), and a Best-of-N (BoN) sampling method. Their results show that the benchmark methods (Averaging and Voting) reliably boost performance, while BoN sampling does not transfer well from mathematical domains.
@@
PromotionGo (Huang,Ziyi et al., 2025) submitted only to the MP-Task A with an XLM-R–based system, ranking first. They deployed three main strategies to develop a competitive system: data augmentation, including lexical swaps, prompt-based reformulation, and large-scale back-translation into nine languages; optimization for alignment to the evaluation metric (Manhattan Distance) by using L1 loss as a loss function; ensemble learning, by training multiple models on shuffled data splits and averaging predictions to improve robustness.
@@
Uncertain Mis(Takes) Staliūnaitė,Ieva and Vlachos,Andreas (2025) addressed only the VEN-Task A, ranking first. They aim to quantify ambiguity in NLI instances, relying on the hypothesis that if a given instance is ambiguous, then the explanations for different labels will not entail one another. For each item, they generate 128 LLM explanations. With a fine-tuned entailment model they cluster them and quantify their Semantic Entropy (SE). The explanation clusters’ SE scores are combined with text embeddings for soft label distribution prediction.
@@
NLP-ResTEAM O. et al. (2025) proposed a multi-task architecture. Special ‘tokens’ are added to the input, including several tokens aiming at modeling the annotators based on their ID, their demographic features, their annotation behavior, or combinations of those. The system produces two outputs from a textual input and an annotator’s information: one is a soft-label, the other a prediction of that specific annotator’s (hard) label.
@@
LPI-RIT Sawkar,Mandira et al. (2025) builds upon the DisCo (Distribution from Context) architecture Weerasooriya et al. (2023), a neural model that jointly predicts item-level, annotator-level, and per-annotator label distributions. They tackled both soft-label and perspectivist tasks simultaneously. They also introduced several extensions to DisCo, such as integrating annotator metadata through pretrained sentence encoders, and modified loss functions to better align with evaluation metrics.
@@ 6.1 Results and statistics
Table 3 and 4 report the overall leaderboard for Task A and Task B respectively. If a team did not submit predictions for a particular dataset or task, we used the random baseline results to compute the overall ranks and average positions. Ranks were calculated with statistical ties taken into account. Specifically, we used the Wilcoxon signed-rank test at the instance level to identify clusters of tied systems. Predictions that were not significantly different (p = 0.05) from the top-performing system in a given cluster were considered ties. A new cluster was formed when a system’s performance was found to be statistically different from that of the best-performing system in the previous cluster
Leadboards for each specific dataset are reported in Appendix B.